Prove that: tan5x+tan3x/tan5x-tan3x=4cos2xcos4x

Grade:12

3 Answers

Latika Leekha
8 years ago
Hello student,
We first consider the L.H.S. and try to reach the R.H.S.
L.H.S = \frac{tan 5x + tan 3x}{tan 5x - tan 3x}
= \frac{\frac{sin 5x}{cos 5x}+\frac{sin 3x}{cos 3x}}{\frac{sin 5x}{cos 5x}-\frac{sin 3x}{cos 3x}}
= \frac{sin 5x cos 3x +sin 3x cos 5x}{sin 5x cos 3x - sin 3x cos 5x}
= \frac{sin (5x+3x)}{sin (5x-3x)}
(Using the formula sin A cos B + cos A sin B = sin (A+B) )
= \frac{sin 8x}{sin 2x}
= 2 sin 4x cos 4x/sin 2x
= 2(2 sin 2x cos 2x)cos 4x/sin 2x
= 4 cos 2x cos 4x.
= R.H.
Sai Sampath
8 years ago
=\frac{sin 5x cos 3x +sin 3x cos 5x}{sin 5x cos 3x - sin 3x cos 5x}\frac{\frac{sin 5x}{cos 5x}+\frac{sin 3x}{cos 3x}}{\frac{sin 5x}{cos 5x}-\frac{sin 3x}{cos 3x}}\frac{tan 5x + tan 3x}{tan 5x - tan 3x}(Using the formula sin A cos B + cos A sin B = sin (A+B)
\frac{sin (5x+3x)}{sin (5x-3x)} 
 = 2 sin 4x cos 4x/sin 2x= 2(2 sin 2x cos 2x)cos 4x/sin 2x= 4 cos 2x cos 4x.\frac{sin 8x}{sin 2x} 
 
 
Om
2 years ago
Hare Krishna Hare Krishna Krishna Krishna Hare Hare
Hare Ram Hare Ram Ram Ram Hare Hare
Hare Krishna Hare Krishna Krishna Krishna Hare Hare
Hare Ram Hare Ram Ram Ram Hare Hare
Hare Krishna Hare Krishna Krishna Krishna Hare Hare
Hare Ram Hare Ram Ram Ram Hare Hare
Hare Krishna Hare Krishna Krishna Krishna Hare Hare
Hare Ram Hare Ram Ram Ram Hare HareHare Krishna Hare Krishna Krishna Krishna Hare Hare
Hare Ram Hare Ram Ram Ram Hare Hare 
nice solution of latika madam